# OpenVIS: Open-vocabulary Video Instance Segmentation

### v2

@@ 5.2 Main Results
MRCNN [17]-BoxTracker $\times$ L-1203 1.4 ${\color[rgb]{.5,.5,.5}0\%}$ 3.0 ${\color[rgb]{.5,.5,.5}0\%}$ 0.9 ${\color[rgb]{.5,.5,.5}0\%}$
MRCNN [17]-STCNTracker [9] $\times$ L-1203 0.9 ${\color[rgb]{.5,.5,.5}0\%}$ 0.7 ${\color[rgb]{.5,.5,.5}0\%}$ 0.6 ${\color[rgb]{.5,.5,.5}0\%}$
MinVIS [22] $\times$ CB-482 1.4 ${\color[rgb]{.5,.5,.5}0\%}$ 5.5 ${\color[rgb]{.5,.5,.5}0\%}$ 0.5 ${\color[rgb]{.5,.5,.5}0\%}$
@@
Detic [45]-SORT [4] L-866 1.9 ${\color[rgb]{.5,.5,.5}15\%}$ 1.8 ${\color[rgb]{.5,.5,.5}0\%}$ 2.5 ${\color[rgb]{.5,.5,.5}18\%}$ 3.4 ${\color[rgb]{.5,.5,.5}100\%}$
Detic [45]-OWTB [28] L-866 2.7 ${\color[rgb]{.5,.5,.5}15\%}$ 2.8 ${\color[rgb]{.5,.5,.5}0\%}$ 1.8 ${\color[rgb]{.5,.5,.5}18\%}$ 4.2 ${\color[rgb]{.5,.5,.5}100\%}$
OV2Seg [32] L-866 3.7 ${\color[rgb]{.5,.5,.5}15\%}$ 3.9 ${\color[rgb]{.5,.5,.5}0\%}$ 2.4 ${\color[rgb]{.5,.5,.5}18\%}$ 11.9 ${\color[rgb]{.5,.5,.5}100\%}$
@@
In this section, we evaluate the performance of the open-world mask proposal, which is a critical component for achieving OpenVIS, using the extensively annotated UVO dataset. As reported in Tab. 2, our mask proposal network, trained solely on the YouTube-VIS dataset with contrastive instance margin loss, outperforming the baseline method [41] traind on YouTube-VIS and even on UVO itself. Compared with the most advanced mask proposal approaches empowered by the Segment Anything Model(SAM) [24] trained with the extensive dataset SA-1B [24], our mask proposal network can also achieve comparable performance.
[…]
MTRCNN [41] YouTubeVIS 7.6 - 9.3
MTRCNN [41] UVO 11.2 - 17.4
TAM [40] SA-1B - 1.7 24.1
SAM-PT [31] SA-1B - 6.7 28.8
InstFormer (Ours) YouTubeVIS 16.7 7.2 24.7
[…]
@@
[…]
Fully-supervised
MaskTrack [41] $\times$ 30.3 51.1 32.6 31.0 35.5
SipMask [6] $\times$ 33.7 54.1 35.8 35.4 40.1
CrossVIS [42] $\times$ 36.3 56.4 38.9 35.6 40.7
VISOLO [16] $\times$ 38.6 56.3 43.7 35.7 42.5
MinVIS [22] $\times$ 47.4 69.0 52.1 45.7 55.7
IDOL [37] $\times$ 49.5 74.0 52.9 47.7 58.1
GenVIS [18] $\times$ 50.0 71.5 54.6 49.5 59.7
DVIS [44] $\times$ 51.2 73.8 57.1 47.2 59.3
MinVIS-CLIP 30.6 51.2 32.0 34.1 40.7
InstFormer (Ours) 51.8 75.6 57.2 49.1 60.0
Open-vocabulary
Detic [45]-SORT [4] 14.6 - - - -
Detic [45]-OWTB [28] 17.9 - - - -
Ov2seg [32] 27.2 - - - -
5.3 Ablation Study
@@
In Tab. 4, comparing Line 1 and Line 2, InstCLIP demonstrates significant improvements over the MinVIS-CLIP baseline, where the masked image is directly input into CLIP for zero-shot classification. Specifically, on BURST, we observe an increase in AP from 2.1 to 3.3 (57% improvement). Similarly, on UVO, the AP rises from 9.0 to 13.1 (46% increase). Notably, for YouTube-VIS, there is a remarkable gain of 18 AP (58% improvement). This enhancement is attributed to InstCLIP’s ability to preserve CLIP’s inherent zero-shot capabilities while being further optimized by the training data from YouTube-VIS. This exceptional characteristic of InstCLIP provides a sound solution for scenarios where both the pursuit of extreme performance for a domain and the ability to handle generalized domains are required.
[…]
@@
Margin $\alpha$1 0 0.5 0.8 1
$AR_{100}$ 21.0 21.1 22.5 22.7 21.9
Table 6: Impact of different similar instance margin $\alpha$ on UVO.
We study the effect of the contrastive instance margin loss to open-world mask proposal on UVO. As shown in Tab. 4 Line 3 and Tab. 6, introducing the contrastive instance margin loss encourages the mask proposal network provide more distinct instances, thereby improving both AP and AR. However, simply pushing all instance queries apart does not achieve this effect. We need to control the hyperparameter margin $\alpha$ to only push instance queries when they are too similar. Empirically, the margin $\alpha$ is set to 0.8.
Instance Tokens for Association.
InstCLIP not only achieves open-vocabulary instance classification effectively and efficiently, but also generates the instance token for each instance. As illustrated in Tab. 4 Line 4 and Tab. 7, this natural byproduct of InstCLIP is beneficial for tracking instances of various categories.
@@
Tab. 7 ablates the key components of the rollout association. Given the instance token provides a richer open-vocabulary instance feature set, the temporal contrastive loss-driven rollout tracker achieves a 10.3% improvement in AP performance in the OpenVIS setting. Additionally, it provides a 3.2% boost in fully-supervised VIS tasks, demonstrating the effectiveness of rollout association in normal fully-supervised tracking instances. This is attributed to the historical information offered by the RNN hidden state, aiding in handling object occlusion and reappearance issues. We also replace the RNN layer with a linear layer or a two-layer MLP with the larger capacity to validate the importance of the historical information provided by the RNN hidden state.
@@
6 Conclusion
In this paper, we proposed a novel OpenVIS model named InstFormer to explore the challenging yet significant computer vision task. Specifically, we first enhance the mask proposal network with the margin instance contrastive loss, improving its capability to generate instance proposals in open-world scenarios. Then, we propose InstCLIP, an enhanced variant of CLIP with instance guidance attention, greatly improving the efficiency and effectiveness of classifying multiple instance proposals within one frame. Meanwhile, the instance token of InstCLIP also benefits instance association across frames. Furthermore, we propose a rollout association mechanism to improve instance association in open-vocabulary scenarios. Last but not least, we propose a comprehensive evaluation benchmark to facilitate further research in this emerging task. Through extensive experiments, we demonstrate the proposed InstFormer achieves state-of-the-art capabilities in OpenVIS and competitive performance in fully supervised VIS.

### v3

@@ Introduction
MRCNN (He et al. 2017)-BoxTracker $\times$ L-1203 1.4 ${\color[rgb]{.5,.5,.5}0\%}$ 3.0 ${\color[rgb]{.5,.5,.5}0\%}$ 0.9 ${\color[rgb]{.5,.5,.5}0\%}$
MRCNN (He et al. 2017)-STCNTracker $\times$ L-1203 0.9 ${\color[rgb]{.5,.5,.5}0\%}$ 0.7 ${\color[rgb]{.5,.5,.5}0\%}$ 0.6 ${\color[rgb]{.5,.5,.5}0\%}$
MinVIS (Huang et al. 2022) $\times$ CB-482 1.4 ${\color[rgb]{.5,.5,.5}0\%}$ 5.5 ${\color[rgb]{.5,.5,.5}0\%}$ 0.5 ${\color[rgb]{.5,.5,.5}0\%}$
@@
Detic (Zhou et al. 2022)-SORT (Bewley et al. 2016) L-866 1.9 ${\color[rgb]{.5,.5,.5}15\%}$ 1.8 ${\color[rgb]{.5,.5,.5}0\%}$ 2.5 ${\color[rgb]{.5,.5,.5}18\%}$ 3.4 ${\color[rgb]{.5,.5,.5}100\%}$
Detic (Zhou et al. 2022)-OWTB (Liu et al. 2022) L-866 2.7 ${\color[rgb]{.5,.5,.5}15\%}$ 2.8 ${\color[rgb]{.5,.5,.5}0\%}$ 1.8 ${\color[rgb]{.5,.5,.5}18\%}$ 4.2 ${\color[rgb]{.5,.5,.5}100\%}$
OV2Seg (Wang et al. 2023) L-866 3.7 ${\color[rgb]{.5,.5,.5}15\%}$ 3.9 ${\color[rgb]{.5,.5,.5}0\%}$ 2.4 ${\color[rgb]{.5,.5,.5}18\%}$ 11.9 ${\color[rgb]{.5,.5,.5}100\%}$
@@
Implementation Details
Model Architecture.
We regard a COCO (Lin et al. 2014)-pretrained Mask2Former (Cheng et al. 2022) as our mask proposal network. By default, the transformer decoder has 100 queries, with a dimension of 256 for the query embedding and transformer decoder. For InstCLIP, we select a ViT-B/32 of CLIP (Radford et al. 2021) as foundation vision transformer. The number of instance tokens of InstCLIP is also set to 100, aligning with the 100 instance queries of the mask proposal network. We initialize the instance tokens and register token using CLIP’s learned class tokens. The text encoder is a 12-layer transformer, the same as that in CLIP. For input prompts, we ensemble 14 prompts (e.g., “a photo of a {category name}”) from (Liang et al. 2022) to boost zero-shot classification ability.
Training.
InstFormer is trained using a two-stage approach and CLIP weights are frozen during the entire training. In first stage, the open-world mask proposal network and InstCLIP (LoRA adapter) are trained for 6k iterations with $\mathcal{L}_{I}$ and instance segmentation loss. Subsequently, we train the rollout tracker in second stage, with all other weights frozen, using $\mathcal{L}_{T}$ for an additional 600 iterations. The whole training is done on 8 V100 GPUs for 3 hours.
Baselines.
To better assess the performance of the proposed InstFormer framework, we introduce several baselines for comparison, as shown in Tab. 1. For fully-supervised methods, we provide STCN Tracker (Athar et al. 2023), Box Tracker (Athar et al. 2023) and MinVIS (Huang et al. 2022). Both the first two methods ultilize the Mask-RCNN (He et al. 2017) but with different tracking strategies as in (Athar et al. 2023). MinVIS (Huang et al. 2022) is an advanced VIS model in fully-supervised VIS task. All of them are trained on full BURST dataset with all 482 categories. For open-vocabulary methods, we employ three approaches: Detic-SORT, Detic-OWTB, and OV2Seg (Wang et al. 2023). The first two methods utilize the open-vocabulary detector Detic (Zhou et al. 2022), paired with the classical multi-object tracker SORT (Bewley et al. 2016) and the state-of-the-art open-world tracker OWTB (Liu et al. 2022), respectively. OV2Seg introduces the CLIP text encoder and a momentum-updated query for tracking, also achieving open-vocabulary video instance segmentation. These baselines may be trained with different datasets. We provide their corresponding training datasets and category numbers in Tab. 1. L-1203 represents the entire LVIS (Gupta, Dollar, and Girshick 2019) dataset with all 1203 categories. L-866 indicates the LVIS subset with 866 frequent categories. B and Y denote BURST (Athar et al. 2023) with 482 categories and YouTube-VIS (Yang, Fan, and Xu 2019) with 40 categories, respectively. C is the COCO (Lin et al. 2014) dataset with 80 categories, which may be used to pretrain some modules. And for clarity, we provide the proportion of novel categories during inference for each approach and setting in the table in gray(higher means more challenging).
Main Results
Training Data AP APc1 $AR_{100}$
MTRCNN (Yang, Fan, and Xu 2019) YouTubeVIS 7.6 - 9.3
MTRCNN (Yang, Fan, and Xu 2019) UVO 11.2 - 17.4
TAM (Yang et al. 2023) SA-1B - 1.7 24.1
SAM-PT (Rajič et al. 2023) SA-1B - 6.7 28.8
InstFormer (Ours) YouTubeVIS 16.7 7.2 24.7
Table 2: Comparison of open-world instance proposal property on UVO. APc1 indicates the class-agnostic AP.
Overall Performance.
We evaluate the overall performance of the InstFormer on the BURST validation set. Since the mask proposal network is pre-trained on COCO and InstFormer is trained on YouTube-VIS, so there are 103 categories have been seen during training. As illustrated in Tab. 1, our proposed InstFormer framework outperforming fully-supervised baselines by a large margin (AP from 1.4 to 4.2). And despite InstFormer seen fewer categories compared to other open-vocabulary baselines, it still achieved state-of-the-art OpenVIS performance (BURST 4.2 AP, LVVISnovel 12.2 AP), demonstrating InstFormer achieves obvious advantages over other methods. Qualitative results can be found in Fig. 1 and Supplementary.
@@
In this section, we evaluate the performance of the open-world mask proposal, which is a critical component for achieving OpenVIS, using the extensively annotated UVO dataset. As reported in Tab. 2, our mask proposal network, trained solely on the YouTube-VIS dataset with contrastive instance margin loss, outperforming the baseline method (Yang, Fan, and Xu 2019) traind on YouTube-VIS and even on UVO itself. Compared with the most advanced mask proposal approaches empowered by the Segment Anything Model(SAM) (Kirillov et al. 2023) trained with the extensive dataset SA-1B (Kirillov et al. 2023), our mask proposal network can also achieve comparable performance.
OV AP AP50 AP75 AR10
Fully-supervised
MaskTrack (Yang, Fan, and Xu 2019) $\times$ 30.3 51.1 32.6 35.5
SipMask (Cao et al. 2020) $\times$ 33.7 54.1 35.8 40.1
CrossVIS (Yang et al. 2021) $\times$ 36.3 56.4 38.9 40.7
VISOLO (Han et al. 2022) $\times$ 38.6 56.3 43.7 42.5
MinVIS (Huang et al. 2022) $\times$ 47.4 69.0 52.1 55.7
IDOL (Wu et al. 2022b) $\times$ 49.5 74.0 52.9 58.1
GenVIS (Heo et al. 2023) $\times$ 50.0 71.5 54.6 59.7
DVIS (Zhang et al. 2023) $\times$ 51.2 73.8 57.1 59.3
MinVIS-CLIP 30.6 51.2 32.0 40.7
InstFormer (Ours) 51.8 75.6 57.2 60.0
Open-vocabulary
Detic-SORT 14.6 - - -
Detic-OWTB 17.9 - - -
Ov2seg 27.2 - - -
@@
Ablation Study
@@
In Tab. 4, comparing Line 1 and Line 2, InstCLIP demonstrates significant improvements over the MinVIS-CLIP baseline, where the masked image is directly input into CLIP for zero-shot classification. Specifically, on BURST, we observe an increase in AP from 2.1 to 3.3 (57% improvement). Similarly, on UVO, the AP rises from 9.0 to 13.1 (46% increase). Notably, for YouTube-VIS, there is a remarkable gain of 18 AP (58% improvement). This shows that InstFormer retains most of CLIP’s zero-shot capabilities while optimizing for YouTubeVIS domain, offering an effective solution for scenarios requiring both extreme domain performance and generalization.
APAll APCom APUncom Once
@@
Instance Tokens for Association.
InstCLIP’s instance tokens not only achieve open-vocabulary classification effectively and efficiently, as illustrated in Tab. 4 Line 4 and Tab. 6, but also aid in tracking instances of any vocabulary.
@@
We study the effect of the contrastive margin loss to open-world mask proposal on UVO. As shown in Tab. 4 Line 3 and Tab. 7 of Supp, introducing the contrastive instance margin loss encourages the mask proposal network provide more distinct instances, thereby improving both AP and AR. More details see Supp.
@@
Tab. 6 ablates the key components of the rollout association. Given the instance token provides a richer open-vocabulary tracking feature, the predictive tracking loss-driven rollout tracker achieves a 10.3% improvement in AP performance in the OpenVIS setting. Additionally, it provides a 3.2% boost in fully-supervised VIS tasks, demonstrating the effectiveness of rollout association in normal fully-supervised tracking instances. RNN offers historical information offered by the hidden state, aiding in handling object occlusion and reappearance issues. We also replace the RNN layer with a linear layer or a two-layer MLP with the larger capacity.
@@
Conclusion
In this paper, we propose InstFormer, a carefully designed framework tailored for the OpenVIS task that achieves powerful open-vocabulary capabilities through lightweight fine-tuning on a limited-category data, to bridge the gap between static, image-level VLMs and the video, instance-level demands of the task. In the InstFormer framework, a mask proposal network encouraged to propose all potential object instances with the margin instance contrastive loss is proposed. Then we propose the InstCLIP to embed open-vocabulary instance tokens efficiently, which enable open-vocabulary instance classification and provide robust open-vocabulary tracking Furthermore, we propose the universal rollout association where the rollout tracker is trained to predict instance tokens of the next frame to achieve universal tracking. Last but not least, we propose a comprehensive evaluation benchmark to facilitate further research in this emerging task. Through extensive experiments, we demonstrate the proposed InstFormer achieves both state-of-the-art capabilities in OpenVIS and competitive performance in fully supervised VIS.
A. Evaluation Benchmark
In our evaluation benchmark, the model is trained on the YouTube-VIS dataset with 40 categories and is tasked with inference on BURST, including an additional 404 unseen categories from the dataset. Given the COCO pretrained mask proposal network, the model has encountered a total of 103 categories during training. There is concurrent work, LV-VIS (Wang et al. 2023), which also aims to establish a benchmark for open-vocabulary video instance segmentation. It trains the model on 866 categories from LVIS datasets and performs inference on LV-VIS with an additional 555 novel categories. While LV-VIS contributes to the community, it’s noteworthy that our setting is more challenging, enabling the evaluation of open-vocabulary capabilities. As illustrated in Fig. 4, our setting involves exposure to a limited number of seen categories (blue) in the training data, while the unseen categories (green) for inference are numerous and distinctly different from the seen categories. In contrast, the LV-VIS setting has seen categories occupying most of the space and being close to the unseen categories.
Figure 4: Comparison of categories distribution space. Category features are extracted using the text encoder of CLIP. The figure is presented after applying the t-SNE dimensionality reduction.
B. Ablation of Contrastive Instance Margin
Introducing the contrastive instance margin loss encourages the mask proposal network to provide more distinct instances. However, simply pushing all instance queries apart does not achieve this effect. It’s crucial to control the hyperparameter margin $\alpha$ so that instance queries are only pushed apart when they are too similar. Empirically, the margin $\alpha$ is set to 0.8.
Margin $\alpha$1 0 0.5 0.8 1
$AR_{100}$ 21.0 21.1 22.5 22.7 21.9
Table 7: Impact of similar instance margin $\alpha$ on UVO.
C. Ablation of Tuning Weight
We also ablate the impact of tuning different pretrained weights of CLIP. As shown in Tab. 8, tuning the query and value projectors in the attention layer leads to a more favorable trade-off performance for both common and uncommon categories.
LoRA APAll APCom APUncom
FFN 3.5 6.5 2.9
Proj. Q,V 3.9 7.4 3.1
Proj. Q,K,V 3.6 6.5 3.0
FNN + Proj. Q,K,V 3.7 5.8 3.3
Table 8: Impact of inserting LoRA into different layers of pre-trained CLIP, on BURST.
D. Qualitative Results
We present qualitative results from the BURST evaluation, where inference involves 482 categories in Fig. 5. The proposed InstFormer successfully recognizes uncommon categories such as “lawn_mower” and even demonstrates the ability to distinguish between “skateboard” and “ski”, as well as “helicopter” and “airplane”.
Figure 5: Qualitative results on BURST. The proposed InstFormer successfully recognizes uncommon categories such as “lawn_mower” and even demonstrates the ability to distinguish between “skateboard” and “ski”, as well as “helicopter” and “airplane”.